How do you solve this expression:
if f(x+5)=f(x)+5, and f(5)=10 then f(50)= ?

Answer by venugopalramana(3286)

How do you solve this expression:
if f(x+5)=f(x)+5, and f(5)=10 then f(50)= ?
put x=5 first
f(5+5)=f(5)+5=10+5
f(10)=15
now put x=10....to get....f(10+5)=f(15)=15+5=20..we can proceed forward like this to get f(50)
similarly we get proceeding backwards..
f(50)=f(45+5)=f(45)+5=f(40)+2*5=f(35)+3*5=f(30)+4*5=f(25)+5*5=f(20)+6*5=f(15)+7*5
=f(10)+8*5=f(5)+9*5=10+45=55